### 31 Mar 2026 v1

# Thomas–Fermi equation revisited: A variation on a theme by Majorana

Abstract.
Majorana found a way to exploit the scaling properties of the Thomas–Fermi equation for converting this second-order differential equation into one of first order. We explore his method for the familiar neutral-atom solution and extend it to the solution that is relevant for weakly ionized atoms. Various integrals and other quantities with importance for atomic physics are recalculated and their values compared with the ones obtained in the 1980s by more tedious numerical procedures.
Key words and phrases:
Thomas–Fermi equation, Majorana’s scaling transformation, homology
2010 Mathematics Subject Classification:
34A05, 35Q99, 53C29, 81V45
(Posted on the arXiv on 31 March 2026.)
1. Introduction
In 1927 and 1928, respectively, Thomas [9] and Fermi [7] managed to combine concepts of the old quantum theory (before mid-1925, that is) with the then-recent methods of quantum mechanics — which grew rapidly from the seminal works of Heisenberg, Schrödinger, and others — for a first successful model of many-electron atoms. The central mathematical object, the second-order differential equation now known as the Thomas–Fermi (TF) equation, immediately caught Majorana’s attention, who had the profound insight that there are scaling properties that can be exploited for establishing an equivalent first-order differential equation.
Most unfortunately, Majorana did not communicate his observations but, fortunately, they were published more than sixty years after his mysterious disappearance, namely when his private notes appeared in print thanks to the commendable effort by Esposito, Majorana Jr., van der Merwe, and Recami [6]. Moreover, Esposito provided a pedagogical exposition of Majorana’s method, supplemented by additional technical details [5].
In this contribution, we further elaborate on Majorana’s first-order equation and use it to compute various numbers with relevance for atomic physics — numbers that were obtained with more tedious methods before; we refer, in particular, to the monograph of 1988 [4] and the works cited therein. In addition to the much studied solution of the TF equation of the original model for neutral atoms, there is another solution that is important for weakly ionized atoms; we apply Majorana’s procedure to that case, too, to find and then explore the corresponding Majorana-type first-order equation.
In Sec. 2, we set the stage by recalling the second-order TF equation and introducing three scale-invariant functions that are constructed from the solutions of the TF equation. Next, in Sec. 3, we establish Majorana’s first-order equation by an argument that is closer in spirit to that in [1] than to the original one in Majorana’s notes, as recalled by Esposito [5]. The analogous argument for the second solution of the TF equation is presented in Sec. 4, where we report the other Majorana-type first-order equation.
Sections 5 and 6 deal with the power series for the two equations, respectively. The first series is the one that Esposito introduced [5] and the second is constructed in close analogy. Both series are used for computing anew the basic numbers associated with the TF functions, which are then compared with the numbers found in the 1980s. Further applications, in particular the computation of various relevant integrals involving the neutral-atom TF function, are the subject matter of Sec. 7. These are then used in Sec. 8 for calculating anew the coefficients in the formulas for the binding energy of neutral atoms and also the leading term in the expression for the ionization energy.
We close with a summary and outlook in Sec. 9.
2. Setting the stage
In the familiar Thomas–Fermi (TF) equation,
[…]
the variable $x$ is positive with x=0{x=0} included for some solutions and the function values $f(x)$ are real. For later reference, we note that
(2.2) $7x^{-1/2}f(x)^{5/2}$ $=\frac{\mathrm{d}}{\mathrm{d}x}\Bigl(5f(x)f^{\prime}(x)-5xf^{\prime}(x)^{2}+4xf(x)f^{\prime\prime}(x)\Biggr)\,$,
$7f^{\prime}(x)^{2}$ $=\frac{\mathrm{d}}{\mathrm{d}x}\Bigl(2f(x)f^{\prime}(x)+5xf^{\prime}(x)^{2}-4xf(x)f^{\prime\prime}(x)\Bigr)\,$,
$12f(x)^{4}-13f^{\prime}(x)^{3}$ $=\frac{\mathrm{d}}{\mathrm{d}x}\Bigl(12xf(x)f^{\prime}(x)f^{\prime\prime}(x)-3f(x)f^{\prime}(x)^{2}-10xf^{\prime}(x)^{3}\Bigr)$
are valid identities for all solutions of (2.1) where $f(x)\geq 0$}.
We are interested in two particular solutions of the TF equation, which are specified by
(2.3) $(\textrm{i})\ 0\leq x<\infty:\quad f(0)$ $=1\,,\quad f^{\prime}(0)=-B\,$,
$\phantom{x\geq 0\,:\quad}f(x)$ $=\frac{144}{x^{3}}\Bigl(1-\beta x^{-\gamma}+\cdots\Bigr)\mbox{\ for\ }x\gg 1\,$,
and
(2.4) $(\textrm{ii})\ 0<x\leq 1:\quad f(1)$ $=0\,,\quad f^{\prime}(1)=-\Lambda^{2}\,$,
$\phantom{0<x\leq 1\,:\quad}f(x)$ $=\frac{144}{x^{3}}\Bigl(1-\alpha x^{\sigma}+\cdots\Bigr)\mbox{\ for\ }0<x\ll 1\,$,
where $\gamma=\bigl(\sqrt{73}-7\bigr)/2$} and $\sigma=\bigl(\sqrt{73}+7\bigr)/2$} and the ellipses stand for higher powers of $x^{-\gamma}$ or $x^{\sigma}$, respectively. We know from previous work (see [4] and the references therein) that [^1] In statements like these, it is understood that the initial digits of the decimal representations of real numbers are given. For the physics applications in Sec. 8, six digits are enough.
[…]
are valid digits. We write $F(x)$ for the solution of case (i), which is relevant for the TF model of neutral atoms, and $\Phi(x)$ for the solution of case (ii), which informs us about weakly ionized atoms. See Fig. 1 for the graphs of the two particular solutions of the TF equation.
Figure 1. The TF functions of the two kinds. Curve a is the graph of $x\mapsto F(x)$} for case (i) and curve b is the graph of $\Phi(x)/\Lambda^{2}\mapsfrom x$} for case (ii).
If $x\mapsto f(x)$ solves the TF equation (2.1), then $x\mapsto f_{\lambda}(x)=\lambda^{3}f(\lambda x)$ with $\lambda>0$ is another solution. Following Majorana’s guidance [5], we consider three scale-invariant combinations, [^2] Or homology-invariant combinations [1], if you prefer this terminology. namely
(2.6) $P(x)$ $=x^{3/2}f(x)^{1/2}\,$, $\mathrm{d}P$ $=\frac{\mathrm{d}x}{x}\biggl(\frac{3}{2}P-\frac{1}{2}P^{-1}Q\biggr)$
$=\frac{\mathrm{d}x}{x}\biggl(\frac{3}{2}P-\frac{1}{2}P^{5/3}R\biggr)\,$,
(2.7) $Q(x)$ $=-x^{4}f^{\prime}(x)\,$, $\mathrm{d}Q$ $=\frac{\mathrm{d}x}{x}\Bigl(4Q-P^{3}\Bigr)$
$=\frac{\mathrm{d}x}{x}\biggl(4Q-Q^{9/8}R^{-9/8}\biggr)\,$,
(2.8) $R(x)$ $=-f(x)^{-4/3}f^{\prime}(x)\,$, $\mathrm{d}R$ $=\frac{\mathrm{d}x}{x}\biggl(\frac{4}{3}R^{2}P^{2/3}-P^{1/3}\Biggr)$
$=\frac{\mathrm{d}x}{x}\biggl(\frac{4}{3}R^{7/4}Q^{1/4}-R^{-1/8}Q^{1/8}\Biggr)$
with $Q^{3}=R^{3}P^{8}$. They are scale-invariant in the sense that the replacement $f\to f_{\lambda}$ yields $P(x)\to P(\lambda x)$, $Q(x)\to Q(\lambda x)$, and $R(x)\to R(\lambda x)$ without additional powers of $\lambda$ as overall factors.
We wish to find one of $P,Q,R$ as a function of another and choose the pair in accordance with the following endpoint values:
[…]
so that we want $R$ as a function of $P$ in case (i) and $Q$ as a function of $P$ in case (ii).
3. Majorana’s equation for case (i)
In case (i), we look for $R$ as a function of $P$. As a consequence of (2.6) and (2.8), we have the differential equation
(3.1) $\frac{\mathrm{d}R}{\mathrm{d}P}=-\frac{2}{3}P^{-2/3}\,\frac{3-4P^{1/3}R^{2}}{3-P^{2/3}R}\,.$
Keeping the values in (2.9) in mind, we put
(3.2) $P=12t^{3}\,,\quad R=\biggl(\frac{3}{16}\biggr)^{1/3}u(t)$
and arrive at Majorana’s equation [5, eq. (27)],
(3.3) $\frac{\mathrm{d}u(t)}{\mathrm{d}t}=-8\frac{1-tu(t)^{2}}{1-t^{2}u(t)}\,$,
with
(3.4) $u(1)=1\,,\quad u(0)=\biggl(\frac{16}{3}\biggr)^{1/3}B\quad\mbox{and}\quad u^{\prime}(1)=-2(1-\gamma)\,;$
see Fig. 2 for the graph of $t\mapsto u(t)$}.
Figure 2. Graphs of the functions $t\mapsto u(t)$} and $s\mapsto v(s)$}.
We then find $x$, $F(x)$, and $F^{\prime}(x)$ in terms of $t$ and $u(t)$ from
(3.5) $\frac{\mathrm{d}x}{x}$ $=\mathrm{d}P\,\biggl(\frac{3}{2}P-\frac{1}{2}P^{5/3}R\biggr)^{-1}$
$=2\frac{\mathrm{d}t}{t}\,\frac{1}{1-t^{2}u(t)}=2\frac{\mathrm{d}t}{t}+2\,\mathrm{d}t\,\frac{tu(t)}{1-t^{2}u(t)}$
so that [5, eqs. (42) and (43)]
(3.6) $x=\lambda t^{2}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$2U(t)$}}\,,\quad F(x)=\frac{144}{\lambda^{3}}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-6U(t)$}}\,,\quad-F^{\prime}(x)=\frac{432}{\lambda^{4}}u(t)\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-8U(t)$}}$
with $\lambda^{3}=144$} and
(3.7) $U(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\,\frac{t^{\prime}u(t^{\prime})}{1-{t^{\prime}}^{2}u(t^{\prime})}$
where the value of the scale parameter $\lambda$ is determined by the requirement of $F(x=0)=1$ in (2.3).
Together, then, we obtain the case-(i) solution of the second-order TF differential equation in (2.1) by solving Majorana’s first-order differential equation in (3.3) and evaluating the integral in (3.7). While the reduction from second to first order is an obvious simplification in itself, there is the additional great benefit that the Majorana function $t\mapsto u(t)$} has a power series that converges in the whole range of $0\leq t\leq 1$} (see [5] and Sec. 5), whereas the two power series for the TF function $x\mapsto F(x)$} have finite, non-overlapping intervals of convergence [4]. [^3] More specifically, the series in powers of $x^{-\gamma}$ converges for $x\gtrsim 30$} and that in powers of $x^{1/2}$ converges for $x\lesssim 0.16$}.
Now, turning to $x\gg 1$, we have
(3.8) $\beta=\biggl(1-\frac{x^{3}F(x)}{144}\biggr)x^{\gamma}\Biggr|_{x\to\infty}$ $=\Bigl(1-t^{6}\Bigr)\Bigl(\lambda t^{2}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$2U(t)$}}\Bigr)^{\gamma}\Biggr|_{1>t\to 1}$
$=6(12)^{2\gamma/3}\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2\gamma\overline{U}$}}$
with
(3.9) $\overline{U}=\int_{0}^{1}\mathrm{d}t\,\biggl(\frac{1}{2\gamma}\frac{1}{1-t}-\frac{tu(t)}{1-t^{2}u(t)}\biggr)\,.$
See Fig. 3 for the graph of $t\mapsto 2\gamma\bigl(U(t)+\overline{U}\bigr)+\log(1-t)$}. [^4] Here and in later occurrences, $\log(\ )$ denotes the natural logarithm.
Figure 3. The sums in (5.4) and (6.4) as functions of $t$ or $s$. Curve a is the graph of $t\mapsto 2\gamma\bigl(U(t)+\overline{U}\bigr)+\log(1-t)$}; curve b is the graph of $s\mapsto(2\sigma/3)\bigl(V(s)+\overline{V}\bigr)+\log(1-s)$}.
4. The analog of Majorana’s equation for case (ii)
In case (ii), we have the differential equation
(4.1) $\frac{\mathrm{d}Q}{\mathrm{d}P}=-2\frac{4PQ-P^{4}}{Q-3P^{2}}$
for finding $Q$ as a function of $P$. We put
(4.2) $P=12s\,,\quad Q=432v(s)$
and obtain
(4.3) $\frac{\mathrm{d}v(s)}{\mathrm{d}s}=-\frac{8}{3}\frac{sv(s)-s^{4}}{v(s)-s^{2}}$
with
(4.4) $v(1)=1\,,\quad v(0)=\frac{1}{432}\Lambda^{2}\quad\mbox{and}\quad v^{\prime}(1)=-\frac{2}{3}(\sigma-3)\,;$
see Fig. 2 for the graph of $s\mapsto v(s)$.
We then find $x$, $\Phi(x)$, and $\Phi^{\prime}(x)$ in terms of $s$ and $v(s)$ from
(4.5) $\frac{\mathrm{d}x}{x}=\mathrm{d}P\,\biggl(\frac{3}{2}P-\frac{1}{2}P^{-1}Q\biggr)^{-1}=-\frac{2}{3}\frac{\mathrm{d}s\,s}{v(s)-s^{2}}$
so that
[…]
$-\Phi^{\prime}(x)$ $=432v(s)\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$(8/3)V(s)$}}$
with
(4.7) $V(s)=\int_{0}^{s}\mathrm{d}s^{\prime}\,\frac{s^{\prime}}{v(s^{\prime})-{s^{\prime}}^{2}}\,$,
where no additional scale factor is needed to ensure that $x\to 1$ when $s\to 0$.
The remarks about case (i) in the paragraph between (3.7) and (3.8) are equally valid for case (ii): The second-order TF equation is converted into the first-order equation in (4.3) and the integration in (4.7), whereby $s\to v(s)$} has a power series that converges for all $s$ in the range $0\leq s\leq 1$} (see Sec. 6), whereas the two power series for $x\to\Phi(x)$} have finite, yet overlapping, intervals of convergence [4]. [^5] More specifically, the series in powers of $x^{\sigma}$ converges for $x\lesssim 0.94$} and that in powers of $(1-x)^{1/2}$ converges for $x\gtrsim 0.88$}.
Now, turning to $0<x\ll 1$, we have
(4.8) $\alpha=\biggl(1-\frac{x^{3}\Phi(x)}{144}\biggr)x^{-\sigma}\Biggr|_{0<x\to 0}$ $=\Bigl(1-s^{2}\Bigr)\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$(2\sigma/3)V(s)$}}\Biggr|_{1>s\to 1}$
$=2\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-(2\sigma/3)\overline{V}$}}$
with
(4.9) $\overline{V}=\int_{0}^{1}\mathrm{d}s\,\biggl(\frac{3}{2\sigma}\frac{1}{1-s}-\frac{s}{v(s)-s^{2}}\biggr)\,.$
See Fig. 3 for the graph of $s\mapsto(2\sigma/3)\bigl(V(s)+\overline{V}\bigr)+\log(1-s)$}.
5. Numerical procedure for case (i)
Esposito [5] notes that an expansion of $u(t)$ in powers of 1t1-t is useful,
(5.1) $u(t)$ $=\sum_{n=0}^{\infty}a_{n}(1-t)^{n}$
[…]
and reports the recurrence relation for the $a_{n}$s, which is equivalent to
(5.2) $a_{n}$ $=\Bigl(2n+16-(n+1)a_{1}\Bigr)^{-1}$
$\mathrel{\phantom{=}}{}\times\Biggl(\frac{1}{2}(n+15)\sum_{m=0}^{n-1}a_{m}a_{n-1-m}-(n+8)\sum_{m=1}^{n-1}a_{m}a_{n-m}$
$\mathrel{\phantom{=}}{}\hphantom{\times\Biggl(}\ +\frac{1}{2}(n+1)\sum_{m=2}^{n-1}a_{m}a_{n-1-m}\Biggr)$
for n>1{n>1}, where the third sum is empty for n=2{n=2}, when
(5.3) $a_{2}=\frac{17-10a_{1}}{20-3a_{1}}a_{1}=\frac{6497-755\sqrt{73}}{152}\,.$
For their values, see [5, Table 1] or the entries in Table 1 and the lin-log graph in Fig. 4; the graph of $u(t)$ in Fig. 2 results from this expansion.
Table 1. The first 2020 coefficients in the power series for $u(t)$ in (5.1), $U(t)$ in (5.4), $v(s)$ in (6.1), and $V(s)$ in (6.4).
$n$ $a_{n}$ $\tilde{a}_{n}$ $b_{n}$ $\tilde{b}_{n}$
1 0.4559 9625 0.2898 1876 5.1813 3458 0.4290 2657
2 0.3044 5508 0.0901 7378 -2.9584 0435 0.1483 3053
3 0.2221 7976 0.0409 2810 0.2678 6614 0.0492 0658
4 0.1682 1262 0.0221 2613 0.0021 4562 0.0168 5127
5 0.1298 0407 0.0132 4507 -0.0070 8457 0.0061 1999
6 0.1013 0020 0.0084 6769 -0.0037 2358 0.0023 9137
7 0.0796 3516 0.0056 6326 -0.0015 1117 0.0010 0172
8 0.0629 2304 0.0039 1218 -0.0005 7357 0.0004 4236
9 0.0499 0531 0.0027 6832 -0.0002 2091 0.0002 0213
10 0.0396 9618 0.0019 9536 -0.0000 9025 0.0000 9430
11 0.0316 4979 0.0014 5926 -0.0000 3946 0.0000 4458
12 0.0252 8385 0.0010 7976 -0.0000 1817 0.0000 2129
13 0.0202 3221 0.0008 0668 -0.0000 0862 0.0000 1026
14 0.0162 1359 0.0006 0754 -0.0000 0415 0.0000 0498
15 0.0130 1011 0.0004 6071 -0.0000 0201 0.0000 0244
16 0.0104 5183 0.0003 5144 -0.0000 0098 0.0000 0120
17 0.0084 0559 0.0002 6946 -0.0000 0048 0.0000 0060
18 0.0067 6661 0.0002 0755 -0.0000 0024 0.0000 0030
19 0.0054 5216 0.0001 6051 -0.0000 0012 0.0000 0015
20 0.0043 9678 0.0001 2458 -0.0000 0006 0.0000 0007
Figure 4. Lin-log plot of the coefficients in (5.1), (5.4), (6.1), and (6.4) for $1\leq n\leq 200$} (or $5\leq n\leq 204$ for $b_{n}$). The large-$n$ slopes are those of geometric sequences with convergence for $\mathopen{\boldsymbol{|}}1-t\mathclose{\boldsymbol{|}}\lesssim 1.20$} and $\mathopen{\boldsymbol{|}}1-s\mathclose{\boldsymbol{|}}\lesssim 1.84$}, respectively, for the corresponding power series.
There is a related power series for the integrand in (3.7) and (3.9),
(5.4) $\frac{\mathrm{d}}{\mathrm{d}t}U(t)$ $=\frac{tu(t)}{1-t^{2}u(t)}=\frac{1}{2\gamma}\frac{1}{1-t}{\left(1-\sum_{n=1}^{\infty}n\tilde{a}_{n}(1-t)^{n}\right)}\,$,
$U(t)$ $=\frac{1}{2\gamma}{\left(-\log(1-t)+\sum_{n=1}^{\infty}\tilde{a}_{n}(1-t)^{n}-\sum_{n=1}^{\infty}\tilde{a}_{n}\right)}\,$,
with
(5.5) $n\tilde{a}_{n}=a_{n}^{\prime}-a_{n}^{\prime\prime}+\sum_{m=1}^{n-1}m\tilde{a}_{m}a^{\prime\prime}_{n-m}\,$,
where
(5.6) $a_{1}^{\prime}=1-a_{1}\,,\quad a_{1}^{\prime\prime}=\bigl(a_{2}-2a_{1}\bigr)/(2\gamma)$
and
(5.7) $a^{\prime}_{k}=a_{k-1}-a_{k}\,,\quad a_{k}^{\prime\prime}=\bigl(a_{k+1}-2a_{k}+a_{k-1}\bigr)/(2\gamma)\quad\mbox{for}\quad k>1\,;$
see Table 1 and Fig. 4. The latter informs us that both series converge for all $t$ in the whole range of $0\leq t\leq 1$ so that [5, eq. (39)]
(5.8) $\biggl(\frac{16}{3}\biggr)^{1/3}B=u(0)=\sum_{n=0}^{\infty}a_{n}$
and
(5.9) $\log\bigl(6(12)^{2\gamma/3}/\beta\bigr)=2\gamma\overline{U}=\sum_{n=1}^{\infty}\tilde{a}_{n}\,.$
All the digits in (2.5) are confirmed in this way by the algorithm described in this section, which is much easier to implement than the one that yielded the values in the 1980s.
Obviously, each coefficient $a_{n}$ is the sum of a rational number and a rational multiple of $\sqrt{73}$, and so is every coefficient $\tilde{a}_{n}$. Therefore, one can obtain exact expressions for the $a_{n}$s and the $\tilde{a}_{n}$s by writing the rational ingredients as ratios of integers, as exemplified by $a_{0}$, $a_{1}$, $a_{2}$ above. Indeed, Moses computed the first 5,000 digits of $B$ in this way [8]. We are less ambitious and content with the decimal approximations obtained with double precision arithmetic, for which $1+2^{-52}>1+2^{-53}=1$}, so that single-step rounding-off errors affect the 16th digit, inasmuch as $2^{-53}=1.11\times 10^{-16}$.
6. Numerical procedure for case (ii)
The analogous power series for $v(s)$ is
(6.1) $v(s)$ $=s^{2}+\sum_{n=1}^{\infty}b_{n}(1-s)^{n}$
$\quad\mbox{with}\quad b_{1}$ $=2-v^{\prime}(1)=\frac{2}{3}\sigma=\frac{1}{3}\bigl(\sqrt{73}+7\bigr)$
and [^6] Here, $\delta_{a,b}$ is the Kronecker delta symbol.
(6.2) $b_{n}$ $=-\biggl((n+1)b_{1}-\frac{14}{3}\biggr)^{-1}$
$\mathrel{\phantom{=-}}\times{\left(\frac{14}{3}b_{n-1}+8\delta_{n,2}-8\delta_{n,3}+\frac{8}{3}\delta_{n,4}+\frac{1}{2}(n+1)\sum_{m=2}^{n-1}b_{m}b_{n+1-m}\right)}$
for n>1{n>1} with an empty sum for n=2{n=2}, when
(6.3) $b_{2}=-\frac{14b_{1}+24}{9b_{1}-14}=-\frac{469+103\sqrt{73}}{456}\,.$
See Table 1 and Fig. 4; the graph of $v(s)$ in Fig. 2 results from this expansion. The corresponding series for the integrand in (4.7) and (4.9) is
(6.4) $\frac{\mathrm{d}}{\mathrm{d}s}V(s)$ $=\frac{s}{v(s)-s^{2}}=\frac{3}{2\sigma}\frac{1}{1-s}{\left(1-\sum_{n=1}^{\infty}n\tilde{b}_{n}(1-s)^{n}\right)}\,$,
$V(s)$ $=\frac{3}{2\sigma}{\left(-\log(1-s)+\sum_{n=1}^{\infty}\tilde{b}_{n}(1-s)^{n}-\sum_{n=1}^{\infty}\tilde{b}_{n}\right)}$
with
(6.5) $\tilde{b}_{1}$ $=1+b_{2}/b_{1}$
and $\and n\tilde{b}_{n}$ $=b_{n+1}/b_{1}-\sum_{m=1}^{n-1}m\tilde{b}_{m}b_{n+1-m}/b_{1}\quad\mbox{for}\quad n>1\,;$
see Table 1 and Fig. 4. Here, too, we have convergence for the whole range of $0\leq s\leq 1$} so that
(6.6) $\frac{1}{432}\Lambda^{2}=v(0)=\sum_{n=1}^{\infty}b_{n}$
and
(6.7) $\log(2/\alpha)=\frac{2\sigma}{3}\overline{V}=\sum_{n=1}^{\infty}\tilde{b}_{n}\,$,
with all the digits in (2.5) confirmed.
7. Integrals involving the TF function and other matters
7.1. Case (i)
The locus and value of the maximum of $xF(x)$,
(7.1) $\mathop{\mathrm{Max}}_{x}{\left\{xF(x)\right\}}=x_{0}F(x_{0})=0.486\,348\,5380$
[…]
follow after finding $t_{0}=0.496\,342\,166\,063$ as the solution of $3t^{2}u(t)=1$}. A simple search algorithm is repeated bisection of an interval that contains $t_{0}$, perhaps starting with $0<t_{0}<1$}.
Integrals of products of powers of $x$ and powers of $F(x)$ can be evaluated in accordance with
(7.2) $\int_{0}^{\infty}\mathrm{d}x\,x^{k}F(x)^{l}=\frac{6l\,(12)^{2(k+1)/3}}{3l-k-1}\int_{0}^{1}\mathrm{d}t\,t^{2k+1}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2(3l-k-1)U(t)$}}$
for 3l>k+1>0{3l>k+1>0}. The conversion of the $x$ integral to the $t$ integral for the parameterization of (3.6) and (3.7) employs
(7.3) $\mathrel{\phantom{=}}\mathrm{d}\biggl(\frac{x^{k-3l+1}}{k-3l+1}\bigl(x^{3}F(x)\bigr)^{l}\biggr)$
$=\mathrm{d}x\,x^{k}F(x)^{l}-\frac{x^{-(3l-k-1)}}{3l-k-1}\mathrm{d}\Bigl(\bigl(x^{3}F(x)\bigr)^{l}\Bigr)$
$=\mathrm{d}x\,x^{k}F(x)^{l}-\mathrm{d}t\,\frac{x^{-(3l-k-1)}}{3l-k-1}\frac{6l}{t}\bigl(x^{3}F(x)\bigr)^{l}$
$=\mathrm{d}x\,x^{k}F(x)^{l}-\mathrm{d}t\,\frac{6l\,(12)^{2(k+1)/3}}{3l-k-1}\,t^{2k+1}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2(3l-k-1)U(t)$}}\,$,
where $x^{3}F(x)=144t^{6}$} is taken into account; the boundary terms of the integration by parts do not contribute.
For benchmarking, we compute the $t$ interal in
(7.4) $u(0)=\biggl(\frac{16}{3}\biggr)^{1/3}B$ $=\frac{7}{5}\biggl(\frac{16}{3}\biggr)^{1/3}\int_{0}^{\infty}\mathrm{d}x\,x^{-1/2}F(x)^{5/2}$
$=12\int_{0}^{1}\mathrm{d}t\;\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-14U(t)$}}\,$,
which follows from the first or the second identity in (2.2). We find
(7.5) $u(0)$ $=2.774\,615\,643\,934\,192\quad\mbox{from this $t$ integral,}$
[…]
and infer that such $t$ integrals yield thirteen correct digits. Other tests, which exploit
(7.6) $1=\bigl(xF^{\prime}(x)-F(x)\bigr)\Bigr|_{x=0}^{\infty}$ $=\int_{0}^{\infty}\mathrm{d}x\,x^{1/2}F(x)^{3/2}$
$=36\int_{0}^{1}\mathrm{d}t\,t^{2}\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-6U(t)$}}$
or
(7.7) $u(0)=\biggl(\frac{16}{3}\biggr)^{1/3}B$ $=\biggl(\frac{16}{3}\biggr)^{1/3}\int_{0}^{\infty}\mathrm{d}x\,x^{-1/2}F(x)^{3/2}$
$=9\int_{0}^{1}\mathrm{d}t\;\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-8U(t)$}}\,$,
confirm this observation.
The exponential functions of multiples of $U(t)$ in (7.2) and (7.4) to (7.7) are handled with the aid of
(7.8) $\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2\kappa U(t)$}}$ $=\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$2\kappa\overline{U}$}}(1-t)^{\kappa/\gamma}{\left(1-\sum_{n=1}^{\infty}\overline{a}_{n}(\kappa/\gamma)(1-t)^{n}\right)}$
$=(6/\beta)^{\kappa/\gamma}(12)^{2\kappa/3}(1-t)^{\kappa/\gamma}{\left(1-\sum_{n=1}^{\infty}\overline{a}_{n}(\kappa/\gamma)(1-t)^{n}\right)}$
$\quad\mbox{with}\quad\overline{a}_{1}(\mu)$ $=\mu\tilde{a}_{1}$
and $\and\overline{a}_{n}(\mu)$ $=\mu\tilde{a}_{n}-\frac{1}{n}\sum_{m=1}^{n-1}\overline{a}_{m}(\mu)(n-m)\mu\tilde{a}_{m}\quad\mbox{for}\quad n>1\,.$
For the $\kappa$ values we need most, the first ten $\overline{a}_{n}(\kappa/\gamma)$ values are given in Table 2. Table 3 lists values of the integrals of $t^{2k+1}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2(3l-k-1)U(t)$}}$, and the resulting values of the various integrals of products of powers of $x$ and $F(x)$ are reported in Table 4; the last row refers to the integral of $\bigl(-F^{\prime}(x)\bigr)^{3}$, which is related to that of $F(x)^{4}$ by the third identity in (2.2), so that
(7.9) $\int_{0}^{\infty}\mathrm{d}x\,\Bigl(-F^{\prime}(x)\Bigr)^{3}=\frac{3}{13}B^{2}-\frac{12}{13}\int_{0}^{\infty}\mathrm{d}x\,F(x)^{4}\,.$
Table 2. The first ten coefficients $\overline{a}_{n}(\kappa/\gamma)$ for $\kappa=1,2,5,7,\text{\ or\ }11$}.
$n$ $\kappa=1$ $\kappa=2$ $\kappa=5$ $\kappa=7$ $\kappa=11$
1 0.3754 1199 0.7508 2398 1.8770 5995 2.6278 8392 4.1295 3188
2 0.0463 3805 -0.0482 5807 -1.1776 5138 -2.6352 5107 -7.2416 6037
3 0.0179 8357 0.0011 7541 0.2710 8440 1.2470 5199 7.0141 5650
4 0.0093 3964 0.0030 2958 -0.0131 9263 -0.2727 3723 -4.0799 4849
5 0.0055 3367 0.0023 8827 -0.0004 8152 0.0205 9164 1.4458 5545
6 0.0035 2913 0.0017 1448 -0.0000 4477 0.0001 6081 -0.2997 1995
7 0.0023 5782 0.0012 1712 -0.0000 0506 0.0000 1105 0.0319 4957
8 0.0016 2628 0.0008 6893 0.0000 0077 -0.0000 0111 -0.0011 4625
9 0.0011 4805 0.0006 2623 0.0000 0211 -0.0000 0192 -0.0000 1360
10 0.0008 2490 0.0004 5576 0.0000 0247 -0.0000 0152 -0.0000 0071
7.2. Case (ii)
A particular integral of $\Phi(x)$ is relevant, namely
(7.10) $\mathrel{\phantom{=}}-\frac{1}{4(12)^{4}}\int_{0}^{1}\mathrm{d}x\,\frac{1}{x^{5}}\,\frac{\mathrm{d}}{\mathrm{d}x}\Bigl(x^{6}\Phi(x)^{2}\Bigr)$
$=\int_{0}^{1}\mathrm{d}s\,\frac{s^{3}}{x^{5}}=\int_{0}^{1}\mathrm{d}s\,s^{3}\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$(10/3)V(s)$}}=1.056\,061\,2411\,$,
which is computed in analogy with (7.8) from the ingredients in (6.4), (6.5), and (6.7). The 1980s value is $(5/4)\times 0.844\,849=1.056\,061$; all digits are confirmed.
Table 3. Values of $\int_{0}^{1}\mathrm{d}t\,t^{2k+1}\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-2(3l-k-1)U(t)$}}$} for 3l>k+1>0{3l>k+1>0}.
$l$ k=1/2k=-1/2 $k=0$ $k=1/2$ $k=1$
1/2 0.57014 68753 0.28246 66572
1 0.38752 34122 0.11447 54905 0.06381 79642 0.05577 72153
3/2 0.30829 06271 0.06927 85149 0.02777 77778 0.01614 68088
2 0.26212 56210 0.04892 33930 0.01587 35703 0.00723 68823
5/2 0.23121 79703 0.03750 88969 0.01041 66667 0.00399 58174
3 0.20876 26928 0.03026 40417 0.00743 39529 0.00249 48104
7/2 0.19154 29334 0.02528 25781 0.00561 26138 0.00168 92551
4 0.17782 11892 0.02165 98536 0.00441 21461 0.00121 15438
Table 4. Integrals of the TF function for case (i). The first column shows $\kappa=3l-k-1$}. The third column lists the values computed in the 1980s with the correct digits underlined. The slanted number 3 in the last row marks a typo in [4] that was not noticed before.
integrand in 1980s value now computed
$\kappa$ $\int_{0}^{\infty}\mathrm{d}x\,(\cdots)$ in Ref. [4] with (7.2), (7.8)
2 $F(x)$ 1.800 063 94 1.800 063 9396
5 $F(x)^{2}$ 0.615 434 64 0.615 434 6934
11 $F(x)^{4}$ 0.247 701 2723
1 $xF(x)$ 9.194 9.194 252 0822
1 $\bigl(F(x)/x\bigr)^{1/2}$ 3.915 933 3.915 931 4911
$\bigl(-F^{\prime}(x)\bigr)^{3}$ 0.3533345 6 0.353 345 6501
8. Energy formulas
In atomic units ($1\,\mathrm{Ha}=27.211\,386\,246\,\mathrm{eV}$), the binding energy of a neutral atom with $Z$ electrons is
(8.1) $E=c_{7}Z^{7/3}-\frac{1}{2}Z^{2}+c_{5}Z^{5/3}+\cdots$
with the coefficients
(8.2) $c_{7}$ $=\frac{6}{7}(3\pi/4)^{-2/3}B=\frac{6}{7}(3\pi^{2})^{-1/3}u(0)=0.768\,745\,124\,$,
$c_{5}$ $=\frac{11}{8}(3\pi/4)^{-4/3}\int_{0}^{\infty}\mathrm{d}x\,F(x)^{2}$
$=\frac{264}{5}\bigl(3\pi^{2}\bigr)^{-2/3}\int_{0}^{1}\mathrm{d}t\;t\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-10U(t)$}}=0.269\,900\,170\,.$
The ellipsis in (8.1) stands for other contributions to the binding energy. First, there is an oscillatory term that is dominated by
(8.3) $E_{\textrm{osc}}=c_{4}\bigl\lceil{c_{\textrm{p}}Z^{1/3}}\bigr\rfloor\biggl(\frac{1}{4}-\bigl\lceil{c_{\textrm{p}}Z^{1/3}}\bigr\rfloor^{2}\biggr)Z^{4/3}+\cdots\,$,
where $\lceil y\rfloor$ denotes the difference between $y$ and its nearest integer, and
(8.4) $c_{4}$ $=4(3\pi/4)^{-2/3}\frac{F(x_{0})}{x_{0}}\biggl(1-\frac{1}{2}x_{0}\sqrt{x_{0}F(x_{0})}\biggr)^{-1/2}=0.320\,593\,992\,$,
$c_{\textrm{p}}$ $=(3\pi/4)^{1/3}\sqrt{x_{0}F(x_{0})}=0.927\,991\,901$
for the amplitude and the period. Second, there is a relativistic correction with the leading terms
(8.5) $E_{\textrm{rel}}=Z^{2}(Z\alpha^{\ }_{\textrm{fs}})^{2}\Bigl(c^{\textrm{rel}}_{0}-c^{\textrm{rel}}_{1}Z^{-1/3}+c^{\textrm{rel}}_{2}Z^{-2/3}+\cdots\Bigr)\,$,
where $\alpha^{\ }_{\textrm{fs}}=1/137.035\,999\,084$ is Sommerfeld’s fine structure constant and the coefficients are
(8.6) $c^{\textrm{rel}}_{0}$ $=\frac{5\pi^{2}}{24}-\zeta(3)=0.854\,110\,680\,$,
$c^{\textrm{rel}}_{1}$ $=3(3\pi/4)^{-4/3}{\left(B^{2}-\int_{0}^{\infty}\mathrm{d}x\,\bigl(-F^{\prime}(x)\bigr)^{3}\right)}$
$=\frac{\bigl(3\pi^{2}\bigr)^{-2/3}}{13}{\left(30u(0)^{2}+\frac{(24)^{3}}{11}\int_{0}^{1}\mathrm{d}t\,t\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$-22U(t)$}}\right)}$
$=2.075\,028\,109\,$,
$c^{\textrm{rel}}_{2}$ $=2(3\pi/4)^{-2/3}B=2\bigl(3\pi^{2}\bigr)^{-1/3}u(0)=1.793\,738\,623\,.$
When relativistic contributions are ignored, the first ionization energy for $Z\gg 1$ is
(8.7) $I$ $=\frac{6}{7}\biggl(\frac{3\pi}{4}\Lambda\biggr)^{-2/3}+\frac{11}{10}\biggl(\frac{3\pi}{4}\Lambda\biggr)^{-4/3}\biggl(\frac{144}{\Lambda}\biggr)^{2}\int_{0}^{1}\mathrm{d}s\,s^{3}\,\mathnormal{\,\mathrm{e}}^{\mbox{\footnotesize$(10/3)V(s)$}}$
$=0.047\,310\,0726+0.068\,506\,3874$
$=0.115\,816\,460\mathrel{\widehat{=}}3.151\,526\,\mathrm{eV}\,$,
where the integral in (7.10) enters.
9. Summary and outlook
We supplemented Majorana’s first-order differential equation for the neutral-atom TF function with the analogous equation for the TF function for weakly ionized atoms, and explored both equations with the aid of Esposito’s power series and its analog. This provided a rather simple way of computing various numbers of physical significance, thereby confirming the values found in the 1980s by much more tedious numerical calculations.
The TF equation is a particular case of
(9.1) $f^{\prime\prime}(x)=x^{a}f(x)^{b}\quad\mbox{or}\quad f^{\prime\prime}(x)=-x^{a}f(x)^{b}$
with $f_{\lambda}(x)=\lambda^{(a+2)/(b-1)}f(\lambda x)\,.$
Other nonlinear ($b\neq 1$}) equations of this form can be investigated as well, among them the Lane–Emden equation for polytropic gas balls (negative sign and a+b=1{a+b=1}); see [1] in this context and the references therein.
In the context of the TF equation itself, it remains to be seen how Majorana’s transformation helps in studying the solutions specified by
(9.2) $f(0)=1\,,\quad f(x_{0})=0\,,\quad-x_{0}f^{\prime}(x_{0})=q\quad\mbox{for}\quad 0<q<1\,$,
where the value of the positive $x_{0}$ depends on $q$, the degree of ionization. [^7] The two cases in (2.3) and (2.4) are obtained in the limit of $q\to 0$}, $x_{0}\to\infty$} without and with the scaling for $\lambda=x_{0}$}, respectively. This is unexplored territory.
Dedication
It gives me great pleasure to dedicate these notes to Goong Chen with whom I share a passion in algorithms that exploit scaling properties [3].

### 23 Apr 2026 v2

@@
 34A05, 35Q99, 53C29, 81V45
-(Posted on the arXiv on 31 March 2026.)
+(Posted on the arXiv on 23 April 2026.)
 1. Introduction
-In 1927 and 1928, respectively, Thomas [9] and Fermi [7] managed to combine concepts of the old quantum theory (before mid-1925, that is) with the then-recent methods of quantum mechanics — which grew rapidly from the seminal works of Heisenberg, Schrödinger, and others — for a first successful model of many-electron atoms.
-The central mathematical object, the second-order differential equation now known as the Thomas–Fermi (TF) equation, immediately caught Majorana’s attention, who had the profound insight that there are scaling properties that can be exploited for establishing an equivalent first-order differential equation.
+In 1926 and 1927, respectively, Thomas [13] and Fermi [7] managed to combine concepts of the old quantum theory (before mid-1925, that is) with the then-recent methods of quantum mechanics — which grew rapidly from the seminal works of Heisenberg, Schrödinger, and others — for a first successful model of many-electron atoms.
+The central mathematical object, the second-order differential equation now known as the Thomas–Fermi (TF) equation, caught Majorana’s attention, who had the profound insight that there are scaling properties that can be exploited for establishing an equivalent first-order differential equation.
 Most unfortunately, Majorana did not communicate his observations but, fortunately, they were published more than sixty years after his mysterious disappearance, namely when his private notes appeared in print thanks to the commendable effort by Esposito, Majorana Jr., van der Merwe, and Recami [6].
@@
 In addition to the much studied solution of the TF equation of the original model for neutral atoms, there is another solution that is important for weakly ionized atoms; we apply Majorana’s procedure to that case, too, to find and then explore the corresponding Majorana-type first-order equation.
-In Sec.
-2, we set the stage by recalling the second-order TF equation and introducing three scale-invariant functions that are constructed from the solutions of the TF equation.
+We begin with a brief account of Thomas’s and Fermi’s reasoning in Sec.
+2.
+Then, in Sec.
+3, we set the stage by recalling the second-order TF equation and introducing three scale-invariant functions that are constructed from the solutions of the TF equation.
 Next, in Sec.
-3, we establish Majorana’s first-order equation by an argument that is closer in spirit to that in [1] than to the original one in Majorana’s notes, as recalled by Esposito [5].
+4, we establish Majorana’s first-order equation by an argument that is closer in spirit to that in [1] than to the original one in Majorana’s notes, as recalled by Esposito [5].
 The analogous argument for the second solution of the TF equation is presented in Sec.
-4, where we report the other Majorana-type first-order equation.
-Sections 5 and 6 deal with the power series for the two equations, respectively.
+5, where we report the other Majorana-type first-order equation.
+Sections 6 and 7 deal with the power series for the two equations, respectively.
 The first series is the one that Esposito introduced [5] and the second is constructed in close analogy.
@@
 Further applications, in particular the computation of various relevant integrals involving the neutral-atom TF function, are the subject matter of Sec.
-7.
+8.
 These are then used in Sec.
@@
 We close with a summary and outlook in Sec. 10.
+2. How Thomas and Fermi reasoned
+Thomas and Fermi, independently, wondered about many-electron atoms, which were beyond the reach of the then-new quantum mechanics.
+The Schrödinger wave function for $N$ electrons has $2^{N}$ components, each a complex-valued function of the $3N$ electron coordinates.
+A compact analytical expression for the atomic ground-state wave function was out of the question then, as it is now, and there was no method for obtaining meaningful approximations.
+[^1] Hartree’s first step in this direction [9] was still in the future.
+Thomas and Fermi resorted to modeling the atom by picturing the electrons moving independently under the influence of the force associated with an effective potential energy $V(\vecfont{r})$, a real-valued function of an electron’s position vector $\vecfont{r}$, which they approximated by
+(2.1) $V(\vecfont{r})=-\frac{Ze^{2}}{r}+\frac{Ze^{2}}{r_{\mathrm{ch}}^{\ }}+e^{2}\int(\mathrm{d}\vecfont{r}^{\prime})\,\frac{n(\vecfont{r^{\prime}})}{\mathopen{\boldsymbol{|}}\vecfont{r}-\vecfont{r}^{\prime}\mathclose{\boldsymbol{|}}}\,$,
+where $e$ is the elementary charge, $r=\mathopen{\boldsymbol{|}}\vecfont{r}\mathclose{\boldsymbol{|}}$} is the length of $\vecfont{r}$, $(\mathrm{d}\vecfont{r}^{\prime})$ denotes the euclidean volume element associated with $\vecfont{r}^{\prime}$, and $n(\vecfont{r}^{\prime})$ is the density of electrons at $\vecfont{r}^{\prime}$; the first term is the Coulomb potential of the nucleus with charge $Ze$, the second term is the chemical potential parameterized by the distance $r_{\mathrm{ch}}^{\ }$, and the third term is the electrostatic potential of the charge density $en(\vecfont{r})$ of the electrons.
+Note that, if the wave function were at hand, one would obtain $n(\vecfont{r})$ by integrating the sum of the absolute squares of all components over the coordinates of N1{N-1} electrons.
+The second ingredient is the statistical-physics expression for the electron density in terms of the momentum integral over the region where the single-electron energy $\vecfont{p}^{2}/(2m)+V(\vecfont{r})$ is negative, [^2] As regulated by the chemical potential.
+(2.4) $n(\vecfont{r})$ $=\frac{2}{(2\pi\hbar)^{3}}\int(\mathrm{d}\vecfont{p})\;\Biggl\{\begin{array}[]{cl}1&\displaystyle\text{where}\ \vecfont{p}^{2}+2mV(\vecfont{r})<0\\[2.15277pt] 0&\text{elsewhere}\end{array}\Biggr\}$
+$=\frac{1}{3\pi^{2}\hbar^{3}}\Bigl(-2mV(\vecfont{r})\Bigr)_{+}^{3/2}\,$,
+counting two electrons per phase-space volume $(2\pi\hbar)^{3}$; here, $m$ is the electron mass, $\vecfont{p}$ is the electron’s momentum vector, $2\pi\hbar$ is Planck’s constant, and $\bigl(y\bigr)_{+}=\bigl(y+\mathopen{\boldsymbol{|}}y\mathclose{\boldsymbol{|}}\bigr)/2$} suppresses negative values of the real variable $y$.
+The approximations entering (2.1) and (2.4) are semi-classical in nature; both can be refined by including quantum corrections, which have been studied extensively in the century since Thomas’s and Fermi’s seminal work.
+Upon applying the Laplace differential operator on both sides of (2.1),
+(2.5) $r>0\,:\quad-\boldsymbol{\nabla}^{2}V(\vecfont{r})=4\pi e^{2}n(\vecfont{r})\,$,
+and taking into account that an isolated atom is isotropic by writing
+(2.6) $V(\vecfont{r})=-\frac{Ze^{2}}{r}f(\kappa r)\quad\text{with}\quad\kappa\frac{\hbar^{2}}{me^{2}}=8(6\pi)^{-2/3}Z^{1/3}\,$,
+we arrive at the familiar TF differential equation for $f(x)$ in (3.1) below.
+The combination of physical constants that multiplies $\kappa$ is Bohr’s radius, the natural unit of length in atomic physics, and the numerical factor is approximately $61/54$.
+The first steps taken by Thomas and Fermi triggered a large body of work in the next two decades, summarized in Gombás’s classic of 1949 [8].
+In the systematic development of density functional theory from the mid-1960s onward, the TF model is the basic approximation, and one has learned, in which sense the TF approximation is exact when both $N$ and $Z$ are very large.
+Recent accounts on the state of affairs from the physical and the mathematical perspectives are the chapters by Okun and Burke [11] and Siedentop [12], respectively, in the proceedings of a workshop in 2019.
 3. Setting the stage
-In the familiar Thomas–Fermi (TF) equation,
+In the TF equation,
 […]
@@
 For the physics applications in Sec.
-8, six digits are enough.
+9, six digits are enough.
 […]
@@
 They are scale-invariant in the sense that the replacement $f\to f_{\lambda}$ yields $P(x)\to P(\lambda x)$, $Q(x)\to Q(\lambda x)$, and $R(x)\to R(\lambda x)$ without additional powers of $\lambda$ as overall factors.
+In (3.6) and (3.8), f(x)>0{f(x)>0} is assumed; for $f(x)\leq 0$}, put P(x)=0{P(x)=0} and $R(x)^{-1}=0$}.
 We wish to find one of $P,Q,R$ as a function of another and choose the pair in accordance with the following endpoint values:
@@
 with $\lambda^{3}=144$} and
-(3.7) $U(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\,\frac{t^{\prime}u(t^{\prime})}{1-{t^{\prime}}^{2}u(t^{\prime})}$
+(4.7) $U(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\,\frac{t^{\prime}u(t^{\prime})}{1-{t^{\prime}}^{2}u(t^{\prime})}\,$,
 where the value of the scale parameter $\lambda$ is determined by the requirement of $F(x=0)=1$ in (3.3).
@@
 While the reduction from second to first order is an obvious simplification in itself, there is the additional great benefit that the Majorana function $t\mapsto u(t)$} has a power series that converges in the whole range of $0\leq t\leq 1$} (see [5] and Sec.
-5), whereas the two power series for the TF function $x\mapsto F(x)$} have finite, non-overlapping intervals of convergence [4].
+6), whereas the two power series for the TF function $x\mapsto F(x)$} have finite, non-overlapping intervals of convergence [3].
 [^5] More specifically, the series in powers of $x^{-\gamma}$ converges for $x\gtrsim 30$} and that in powers of $x^{1/2}$ converges for $x\lesssim 0.16$}.
@@
 The remarks about case (i) in the paragraph between (4.7) and (4.8) are equally valid for case (ii): The second-order TF equation is converted into the first-order equation in (5.3) and the integration in (5.7), whereby $s\to v(s)$} has a power series that converges for all $s$ in the range $0\leq s\leq 1$} (see Sec.
-6), whereas the two power series for $x\to\Phi(x)$} have finite, yet overlapping, intervals of convergence [4].
+7), whereas the two power series for $x\to\Phi(x)$} have finite, yet overlapping, intervals of convergence [3].
 [^7] More specifically, the series in powers of $x^{\sigma}$ converges for $x\lesssim 0.94$} and that in powers of $(1-x)^{1/2}$ converges for $x\gtrsim 0.88$}.
@@
 9. Energy formulas
-In atomic units ($1\,\mathrm{Ha}=27.211\,386\,246\,\mathrm{eV}$), the binding energy of a neutral atom with $Z$ electrons is
+In atomic units ($1\,\mathrm{Ha}=27.211\,386\,246\,\mathrm{eV}$), the binding energy of a neutral atom with N=Z{N=Z} electrons is
 (9.1) $E=c_{7}Z^{7/3}-\frac{1}{2}Z^{2}+c_{5}Z^{5/3}+\cdots$
@@
 In the context of the TF equation itself, it remains to be seen how Majorana’s transformation helps in studying the solutions specified by
-(9.2) $f(0)=1\,,\quad f(x_{0})=0\,,\quad-x_{0}f^{\prime}(x_{0})=q\quad\mbox{for}\quad 0<q<1\,$,
-where the value of the positive $x_{0}$ depends on $q$, the degree of ionization.
-[^7] The two cases in (2.3) and (2.4) are obtained in the limit of $q\to 0$}, $x_{0}\to\infty$} without and with the scaling for $\lambda=x_{0}$}, respectively.
+(10.2) $f(0)=1\,,\quad f(x_{\mathrm{ch}}^{\ })=0\,,\quad-x_{\mathrm{ch}}^{\ }f^{\prime}(x_{\mathrm{ch}}^{\ })=q\quad\mbox{for}\quad 0<q<1\,$,
+where the value of the positive $x_{\mathrm{ch}}^{\ }$ depends on q=1N/Z{q=1-N/Z}, the degree of ionization.
+[^9] The relation $x_{\mathrm{ch}}^{\ }=q\kappa r_{\mathrm{ch}}^{\ }$} connects with (2.6).
+The two cases in (3.3) and (3.4) are obtained in the limit of $q\to 0$}, $x_{\mathrm{ch}}^{\ }\to\infty$} without and with the scaling for $\lambda=x_{\mathrm{ch}}^{\ }$}, respectively.
 This is unexplored territory.
@@
 It gives me great pleasure to dedicate these notes to Goong Chen with whom I share a passion in algorithms that exploit scaling properties [2].
+Acknowledgment
+I am extremely grateful for the long-standing support from the Centre for Quantum Technologies, Singapore, where part of this work was done.